# Low-Rank Interconnected Adaptation across Layers

### v2

@@ Abstract
Low-rank adaptation (LoRA) is a powerful parameter-efficient fine-tuning method that utilizes low-rank projectors $A$ and $B$ to learn weight updates $\Delta W$ for adaptation targets $W$. Previous research has shown that LoRA is essentially a gradient compressor, performing random projections on the gradient using a fixed projection matrix $A_{0}$. However, this setup restricts the overall weight update to be low-rank, which limits the adaptation performance. In this paper, we propose low-rank interconnected adaptation across layers (Lily). Specifically, we employ a hierarchical framework where low-dimensional projectors (LPs) retained for downward projection at a particular level, while globally-shared high-dimensional projector (HP) experts perform upward projection across all levels of layers. Lily uniquely connects each LP to all HP experts, therefore the gradient projections are no longer dominated by fixed projection matrices, but rather by selective combinations of all the projectors, thereby breaking the low-rank constraint of LoRA. Furthermore, Lily’s cross-layer connections facilitate the capture of intricate information and dependencies across different layers, thereby enhancing the model’s representational capabilities. Experiments across various modalities, architectures, and model sizes underscore Lily’s great performance and efficiency. Code is available on github https://github.com/yibozhong/lily.
@@ 1 Introduction
For foundation models like Transformers (Vaswani et al., 2017b), fine-tuning on downstream tasks is a typical usage, but full fine-tuning (FFT) of large models like large language models (LLMs) incurs huge computational and storage costs and risks forgetting previously learned knowledge (Biderman et al., 2024). Linear probing, which fine-tunes only the final modules like classification heads, addresses these issues but leads to significant performance degradation since it doesn’t update weights from the backbone. To tackle these challenges, parameter-efficient fine-tuning (PEFT) has received significant attention. In PEFT, a model’s backbone weights are frozen, and lightweight trainable modules are introduced to efficiently learn task-specific knowledge. Among all PEFT methods, Low-rank Adaptation (LoRA (Hu et al., 2021)) is one of the most widely applied techniques, especially in LLMs. LoRA introduces a pair of low-rank projection matrices for each adaptation target, consisting of a downward adapter $A$ and an upward adapter $B$, to approximate $\Delta W$ in FFT. Due to its low-rank nature, LoRA offers significant computational and storage savings, effectively alleviating the burdens of FFT while significantly outperforming linear probing by learning the weight updates for backbone weight.
However, LoRA and many subsequent improvements to the method (Miles et al., 2024), (Zhang et al., 2023), (Zhong et al., 2024) have a limitation: the overall learned weight updates $\Delta W$ are restricted to be low-rank, which limits the model performance during adaptation. Previous work (Hao et al., 2024) has shown that LoRA adapters are secretly gradient compressors. In particular, LoRA updates to the weight can be approximated by performing random projection on the gradient, using a fixed random projection matrix $A_{0}$, where $A_{0}$ is the initial weight for $A$. Since this limitation arises because the projection matrix is always fixed in LoRA, using multiple distinct matrices and their variations to perform the gradient projection could avoid restricting the weight updates to a low-rank subspace. This prompts a question: How can we enable high-rank updates by using various projection matrices for an adaptation target?
Figure 1: Dynamics of LoRA and Lily. $N_{l}$ is used to denote the number of layers in the model while $N_{e}$ denoting the number of high-dimensional projector experts from Lily. R is representing the router from Lily.
@@
In this paper, we propose Low-rank interconnected adaptation across layers (Lily), a novel framework for more expressive and performative PEFT. Specifically, we decouple the downward low-dimensional projector (LP) and its corresponding upward high-dimensional projectors (HP), making them not tightly-bonded. Each LP is connected to all the HPs, and vice versa as shown in Fig. 1. This results in a hierarchical structure where LPs are still retained at a particular level to perform downward projection, while all HPs are now globally shared by all the LPs, performing upward projection. Inspired by self-attention Vaswani et al. (2017a), which calculates the relationship between a token and all tokens and obtains attention scores indicating the strength of their relationship, we selectively connect an LP with the HPs based on layer features. The LP extracts features from the current layer, and based on the extracted features, a data-dependent and selective combination of HPs is performed. This is realized by utilizing a router (Shazeer et al., 2017) that outputs a unique weight distribution for HP experts, depending on the current input feature, thereby exhibiting selectivity.
@@
The gradient projection now uses multiple random projection matrices from all the layers and their selective combinations rather than a fixed matrix, thus effectively avoiding the constraint of low-rank updates from LoRA. Furthermore, Lily enables a more comprehensive information access by allowing adapters at each layer to access information from other layers, promoting an interconnected and dynamic learning process, where the adapters can collaborate, share learned knowledge and model dependencies across layers. Overall, our contributions include:
  • We propose Lily, a novel PEFT framework that incorporates cross-layer connections of the projection matrices, thereby breaking the restriction of low-rank weight updates in LoRA.
@@
  • Lily utilizes routers to selectively connect an LP with multiple HP experts, enabling comprehensive information access and therefore expressive adaptation.
@@
  • Extensive experiments are conducted across various modalities, architectures, and model sizes, highlighting Lily’s great performance and efficiency in diverse scenarios.
@@ 2 Related Work
Parameter Efficient Fine-Tuning. Typical usage of foundation models includes pre-training on large datasets and fine-tuning on various downstream tasks. Parameter-efficient fine-tuning (PEFT) thus emerges as a promising field, aiming to fine-tuning the model efficiently with minimal parameters while maintaining performance and preserving previously learned knowledge, addressing drawbacks posed by conventional fine-tuning techniques like full fine-tuning or linear probing. Current PEFT research can be mainly categorized into two types: 1) adapter-based methods (Hu et al., 2021), (Chen et al., 2022), (Pfeiffer et al., 2020b), (Jie and Deng, 2023) (Houlsby et al., 2019b) and 2) prompt-based methods (Tu et al., 2023b) (Tu et al., 2023a). Adapter-based methods introduce lightweight adapters into the Multi-Head Self-Attention (MHSA) or the Feed-Forward Network (FFN) blocks within the Transformer architecture. On the other hand, prompt-based methods append trainable tokens as prompts to the input sequence fed to certain parts of the model.
@@
Among these various PEFT techniques, low-rank adaptation (LoRA (Hu et al., 2021)) stands out as one of the most well-known methods. LoRA introduces a pair of projection matrices $A$ and $B$ per adaptation target $W$. The low-dimension projector (LP) $A$ projects input $x$ to low-dimension space, and the high-dimension projector (HP) $B$ restores it to its original dimension. Multiplying these projection matrices approximates the weight update $\Delta W$ in FFT. Recent work (Hao et al., 2024) has shown that LoRA adapters are essentially performing random projection to the gradient using a fixed matrix. This restricts the learned weight update to low-rank subspace and thus imitating the model performance. Meanwhile, $A$ and $B$ are tightly coupled, therefore the adaptation process only has information access from current layer, without an understanding of information from other layers, which could be beneficial to modeling dependencies across various layers.
@@
Mixture of Experts. Mixture of Experts (MoE) is an active research area that has garnered significant attention, especially in the field of large language models (LLMs). Conditional computation, where different parts of the network are activated on a per-example basis, has been proposed to enhance model capability without increasing computation (Davis and Arel, 2013) (Bengio et al., 2013) (Eigen et al., 2013) (Almahairi et al., 2016). The sparsely-gated MoE layer is introduced to implement this idea, consisting of numerous sub-networks (Shazeer et al., 2017). A trainable gating network, known as a "router", determines the combination of experts for each example. There are already PEFT methods like MoLORA (Zadouri et al., 2023) and MOLA (Gao et al., 2024a) which apply the MoE design concept to PEFT. However, these methods simply treat the adapters combined in LoRA as a single expert. A concurrent research Wu et al. (2024), utilizes LP and HP sub-spaces as the experts but fails to overcome the limitation discussed in previous section.
@@ 3 Methodology
3.1 Preliminaries
LoRA is a Gradient Projector. Adaptation process of low-rank adaptation (LoRA) can be expressed as follows when setting the rank to $d$, a small hidden-dimension:
$W=W_{0}+AB$ (1)
where $W,W_{0}\in\mathbb{R}^{C_{in}\times C_{out}}$ with $C_{in}$ being the input dimension and $C_{out}$ being the output dimension. $W_{0}$ is the frozen weight of a target module and $W$ is the actual value of the target during adaptation runtime. $A\in\mathbb{R}^{C_{in}\times d}$ and $B\in\mathbb{R}^{d\times C_{out}}$ are low-dimension projectors (LP) and high-dimension projectors (HP), with their product multiplied by a scalar value $s$ (omitted in the equation) to approximate weight update $\Delta W\in\mathbb{R}^{C_{in}\times C_{out}}$. We denote the information from $W_{0}$ as the basic knowledge that the pre-trained model already possesses, while $\Delta W=AB$ represents the learned extra knowledge specific to the current downstream task. We denote the initial value of $A$ and $B$ as $A_{0}$ and $B_{0}$, and their update as $\Delta A$ and $\Delta B$. $B_{0}$ is initialized as 0. The approximation is then described as:
$W$ $=W_{0}+(A_{0}+\Delta A)(B_{0}+\Delta B)$ (2)
$=W_{0}+A_{0}\Delta B+\Delta A\Delta B$ (3)
$\approx W_{0}+A_{0}\Delta B$ (4)
It is proved in Hao et al. (2024) that it is the term $A_{0}\Delta B$ that contributes the most to the weight updates. The forward pass given an input in LoRA is:
$y=xW=xW_{0}+xAB$ (5)
where $x\in\mathbb{R}^{C_{in}}$ and $y\in\mathbb{R}^{C_{out}}$ for simplicity. Therefore, the gradient w.r.t. $B$ is:
$\frac{\partial{\mathcal{L}}}{\partial B}$ $=A^{T}x^{T}\frac{\partial{\mathcal{L}}}{\partial y}$ (6)
$=A^{T}(\nabla_{W}{\mathcal{L}})$ (7)
where $x^{T}\frac{\partial{\mathcal{L}}}{\partial y}\in\mathbb{R}^{C_{in}\times C_{out}}$ is the gradient for the weight matrix, which we denote as $(\nabla_{W}{\mathcal{L}})$. Using the approximation in Eq. 2, we can express the adaptation when only updating $B$ using gradient descent:
$W=W_{0}-\eta\sum_{t=0}^{T}\big{[}A_{0}A_{0}^{T}(\nabla_{W}{\mathcal{L}}_{t})\big{]}$ (8)
where $T$ is the current time-step and $\eta$ is the adopted learning rate. The projection matrix here is $A_{0}$ because $A$ is not updated in the approximation, since its change contribute little to the overall weight updates. We can observe that LoRA is essentially performing random projection to the gradient using a fixed random projection matrix $A_{0}$. However, this setup limit the weight update $\Delta W$ to be low-rank, which also limit the adaptation performance.
Lily is a Better Gradient Projector. Consider simply connecting each $A$ with all the $B$s with weights, the forward pass is then:
$W$ $=W_{0}+\sum_{i=1}^{N_{e}}\Big{[}S_{i}\big{[}(A_{0}+\Delta A)(B_{0}^{i}+\Delta B^{i})\big{]}\Big{]}$ (9)
$=W_{0}+\sum_{i=1}^{N_{e}}\big{[}S_{i}(A_{0}\Delta B^{i}+\Delta A\Delta B^{i})\big{]}$ (10)
$\approx W_{0}+A_{0}\sum_{i=1}^{N_{e}}S_{i}\Delta B^{i}$ (11)
where $N_{e}$ is used to denote the number of $B$s in the model and $S$ is a set of weight scores of current $A$ for all the $B$s, each $S_{i}$ is a scalar value representing the weight for corresponding $B^{i}$.
Thanks to the inter-connectivity, each $B$ is also connected to all the $A$s. The gradient w.r.t an $B$ is now:
$\frac{\partial{\mathcal{L}}}{\partial B}$ $=\sum_{i=1}^{N_{l}}C_{i}{A^{i}}^{T}{x^{i}}^{T}\frac{\partial{\mathcal{L}}}{\partial y^{i}}$ (12)
$=\sum_{i=1}^{N_{l}}C_{i}{A^{i}}^{T}(\nabla_{W_{i}}{\mathcal{L}})$ (13)
where $N_{l}$ is the number of $A$s in the model and we denote gradient for each corresponding weight matrix as $\nabla_{W_{i}}{\mathcal{L}}$. $C$ is a set of weight scores where each $C_{i}$ is a weight score from certain $A_{i}$ for current $B$. The forward pass when only updating $B$ can then be expressed as:
$W=W_{0}-\eta\sum_{t=0}^{T}\Big{[}\sum_{1}^{N_{e}}S_{i}\big{[}\sum_{j=1}^{N_{l}}{C_{i,j}A_{0}A_{0}^{j}}^{T}(\nabla_{W_{j}}{\mathcal{L}}_{t})\big{]}\Big{]}$ (14)
where $C_{i,j}$ is used to denote weight scores from various $A^{j},j\in\{1,2,...,N_{l}\}$ to $B_{i},i\in\{1,2,...,N_{e}\}$. It can be observed that with this setup, the projection to the gradient no longer use a fixed $A_{0}$, but rather utilizing all the projection matrix $A_{0}^{i},i\in\{1,2,...,N_{l}\}$ and their overall complex combinations. Note that all the ${A_{0}}^{i}$s are distinct random projection matrices and each weight score ($S_{i}$ and $C_{i,j}$) is unique, data-dependent and thus varies across time-steps. Therefore, the weight updates are not restricted to be low-rank, but rather high-rank. This setup is precisely what Lily adopts. To differentiate from LoRA, we use different notations for the projectors. Specifically, we decouple HPs from their original layers, allowing them to form a model-wide shared module of HP experts ($B$ in LoRA), denoted as $P_{H}^{i}$, where $i\in\{1,2,...,N_{e}\}$ and $N_{e}$ is the number of HP experts. Meanwhile, we fix LPs ($A$ in LoRA) $P_{L}^{i},i\in\{1,2,...,N_{l}\}$ in each layer for downward projection, where $N_{l}$ is the number of LPs. This hierarchical structure makes the projection matrices all interconnected across layers, enabling the modeling of complex dependencies and high-rank updates to the model weights. Detailed model structure and design intuitions are discussed in Appendix A.
3.2 Dynamics of Lily
@@
Downward Projection and Selective Weight Allocation. The process is illustrated in the right half of Fig. 1. Initially, we use an LP to project the input $x\in\mathbb{R}^{N\times C_{in}}$ into its low-dimensional representation $x^{\prime}\in\mathbb{R}^{N\times d}$ where $N$ is the sequence length:
$x^{\prime}=xP_{L}$ (15)
@@
Inspired by the Mixture of Experts (MoE) paradigm, we employ a router $R\in\mathbb{R}^{N_{e}\times d}$ to selectively assign weights to all HP experts based on their relationship to the current layer’s features ($x^{\prime}$). The weight set $S$ is obtained as:
$S=softmax(\sum_{i=1}^{N}(x^{\prime}R^{T})_{i})$ (16)
The router selectively combines experts based on the current layer’s features, enabling smart information integration. For shallower inputs, the router increases attention for experts specializing in shallow-layer knowledge, while deeper inputs favor experts learning deep-layer knowledge.
@@
Weighted Combination of Experts and Upward Projection. Once we obtain the low-dimensional input $x^{\prime}$ , we combine information from all layers using the model-wide shared global HP module. One intuitive approach is to feed $x^{\prime}$ into each HP expert and combine their outputs to obtain the extra knowledge $x_{\Delta}\in\mathbb{R}^{N\times C_{out}}$:
$x_{\Delta}=\sum_{i=1}^{N_{e}}S_{i}\cdot(x^{\prime}P_{H}^{i})$ (17)
where $S\in\mathbb{R}^{N_{e}}$ is the set of weight scores for HP experts, obtained through selective weight allocation. To address efficiency concerns discussed in Appendix A.2, we propose an alternative implementation that is mathematically equivalent and significantly reduces computational burden, described as:
$x_{\Delta}=x^{\prime}(\sum_{i=1}^{N_{e}}S_{i}\cdot P_{H}^{i})$ (18)
@@
Since each $S_{i}$ is a scalar value, the calculation in Eq. 18 is mathematically equivalent to that in Eq. 17, but with significantly improved efficiency. Therefore, the whole computation flow, with input $x\in\mathbb{R}^{N\times C_{in}}$ and output $y\in\mathbb{R}^{N\times C_{out}}$, for an adaptation target module is:
@@
$=xW_{0}+s\cdot xP_{L}(\sum_{i=1}^{N_{e}}(softmax(\sum_{j=1}^{N}(xP_{L}R^{T})_{j}))_{i}\cdot P_{H}^{i})$ (20)
By selectively allocating weights and combining HP experts, Lily enables access to all levels of information during adaptation. Each layer’s target adaptation modules could consider the status and knowledge from all other layers, resulting in a more expressive and comprehensive adaptation. Meanwhile, thanks to its inter-connectivity and selectivity, Lily break the low-rank update constraint of LoRA and enable high-rank updates, as discussed in preliminaries.

### v3

@@ Abstract
[…]
Low-Rank Interconnected Adaptation across Layers
Yibo Zhong1, Jinman Zhao2, Yao Zhou1thanks: Corresponding authorthanks: This work was supported by National Natural Science Foundation of China (Grant 62376172), 1Sichuan University, 2University of Toronto, zhongyibo@stu.scu.edu.cnyaozhou@scu.edu.cn
@@ 1 Introduction
Figure 1: Dynamics of LoRA and Lily. In this 6-layer example with a fixed overall parameter budget, LoRA allocates the same parameter budget to each layer, resulting in small rank updates for the weights. Lily overcomes this by employing a small number of shared adapters with a much larger rank, achieving higher-rank updates while using the same or even a smaller parameter budget. Considering the different characteristics, and to make the adaptation more dynamic, the adapters are mixed according to a data-dependent router, represented by $R$.
@@
Fine-tuning foundation models like Transformers (Vaswani et al., 2017) on downstream tasks is common but costly, especially for large models like LLMs, which incur high computational and storage demands and risk catastrophic forgetting (Biderman et al., 2024). Linear probing alleviates these issues by fine-tuning only the final modules, but suffers from performance loss due to frozen backbone weights. To address this, parameter-efficient fine-tuning (PEFT) freezes the backbone and introduces lightweight modules for task-specific learning. Among PEFT methods, Low-rank Adaptation (LoRA (Hu et al., 2021)) is widely used, particularly for LLMs. LoRA introduces low-rank projection matrices, $A$ and $B$, to approximate weight updates $\Delta W$, achieving significant savings in computation and storage while outperforming linear probing by updating the backbone weights.
However, LoRA and its subsequent improvements (Miles et al., 2024; Zhang et al., 2023; Zhong et al., 2024) face a limitation: the learned weight updates $\Delta W$ are constrained to be low-rank, limiting model performance. A key issue is that LoRA allocates the same parameter budget to each layer, regardless of their importance (Fig. 1). As a result, the rank of each adapter is constrained by the fixed budget, raising the critical question: Can we enable more dynamic, expressive adaptation with high-rank weight updates under the same parameter budget?
@@
In this paper, we propose Low-rank interconnected adaptation across layers (Lily), a novel framework for more expressive and efficient PEFT. Specifically, we decouple $A$ and its corresponding upward $B$, eliminating their tight coupling. Each $A$ is connected to all $B$s, and vice versa, as illustrated in Fig. 1. This creates a hierarchical structure where locally-shared $A$s perform downward projections at specific layers, while globally-shared $B$s perform upward projections across all layers. To enhance dynamism, we selectively connect each $A$ with $B$s based on layer features. $A$ extracts features from the current layer, and a selective mixture of $B$s is performed based on these features, enabled by routers (Shazeer et al., 2017) that generate data-dependent weight distributions for $B$ experts.
@@
The interconnected structure makes the adaptation process more dynamic and flexible, with rich interactions between adapters. By reducing the number of adapters and increasing their rank, Lily achieves higher-rank weight updates than LoRA while using the same or fewer parameters. Additionally, Lily enables comprehensive information access and learning by allowing adapters at each layer to collaborate, share knowledge, and model dependencies across layers. Our key contributions include:
  • We propose Lily, a novel PEFT framework that introduces interconnected adapters, effectively overcoming the limitations of low-rank weight updates in LoRA under the same parameter constraints.
@@
  • Lily utilizes routers to dynamically select and connect an adapter $A$ with multiple adapter $B$ experts, enabling richer information flow and more expressive adaptation dynamics.
@@
  • Extensive experiments are conducted across diverse modalities, architectures, and model scales, demonstrating Lily’s superior performance and efficiency in a wide range of scenarios.
@@ 2 Related Work
Parameter Efficient Fine-Tuning Foundation models are typically pre-trained on large datasets and fine-tuned on downstream tasks. Parameter-efficient fine-tuning (PEFT) seeks to fine-tune models efficiently with minimal parameters while maintaining performance and preserving learned knowledge. It effectively addresses limitations of conventional fine-tuning techniques, like full fine-tuning or linear probing. Current PEFT approaches can be divided into two categories: 1) adapter-based methods (Hu et al., 2021; Chen et al., 2022; Pfeiffer et al., 2020a; Jie and Deng, 2023; Houlsby et al., 2019a; Zhong and Zhou, 2025) and 2) prompt-based methods (Tu et al., 2023a, b). Adapter-based methods insert lightweight adapters into the Multi-Head Self-Attention (MHSA) or Feed-Forward Network (FFN) blocks of the Transformer architecture, while prompt-based methods add trainable tokens to the input sequence.
@@
Among these, low-rank adaptation (LoRA (Hu et al., 2021)) is a well-known technique. It introduces projection matrices $A$ and $B$ for each adaptation target $W$, where $A$ projects input $x$ to a low-dimensional space and $B$ restores it to the original dimension. The product of these matrices approximates the weight update $\Delta W$ in full fine-tuning (FFT). However, this limits the update to a low-rank subspace, which may affect performance. Additionally, $A$ and $B$ are tightly coupled, restricting the adaptation process to information from the current layer, which may hinder the modeling of dependencies across layers.
@@
Mixture of Experts Mixture of Experts (MoE) is an active research area that has received significant attention, especially in the field of large language models (LLMs). Conditional computation, which activates different parts of the network on a per-example basis, has been proposed to enhance model capability without increasing computation (Davis and Arel, 2013; Bengio et al., 2013; Eigen et al., 2013; Almahairi et al., 2016) . The sparsely-gated MoE layer is introduced to implement this idea, consisting of numerous sub-networks (Shazeer et al., 2017). A trainable gating network (router) determines the combination of experts for each example. There are already PEFT methods like MoLORA (Zadouri et al., 2023) and MOLA (Gao et al., 2024a) that apply the MoE design concept to PEFT. However, these methods simply treat the adapters $A$ and $B$ combined in LoRA as a single expert. Concurrent research Wu et al. (2024) utilizes $A$ and $B$ sub-spaces as the experts but fails to overcome the limitation discussed in the previous section. Another concurrent work, HydraLoRA Tian et al. (2024), explores an asymmetric design for LoRA. Unlike our work, we consider the interconnection across layers and deploy a model-wide asymmetric design to enable cross-layer connections. This enables the use of adapters of higher rank than a typical LoRA setup while using the same or fewer overall parameters.
@@ 3 Methodology
3.1 Downward Projection and Selective Weight Allocation
@@
The process is illustrated in the right half of Fig. 1. Initially, we use an $A$ to project the input $x\in\mathbb{R}^{N\times C_{\text{in}}}$ into its low-dimensional representation $x^{\prime}\in\mathbb{R}^{N\times d}$, where $N$ is the sequence length:
$x^{\prime}=xA$ (1)
@@
To enable more parameter efficiency, the number of $A$s can be set to less than the number of layers in the model by sharing the same $A$ across neighboring layers, as illustrated in Fig. 1 and discussed in A. Inspired by the Mixture of Experts (MoE) paradigm, we employ a router $R\in\mathbb{R}^{N_{e}\times d}$ to selectively assign weights to all $B$ experts based on their relationship to the current layer’s features ($x^{\prime}$), where $N_{e}$ represents the number of $B$ experts. A weight set $S\in\mathbb{R}^{N_{e}}$ is obtained as:
$S=\text{softmax}\left(\sum_{i=1}^{N}(x^{\prime}R^{T})_{i}\right)$ (2)
The router selectively mixes the experts based on this data-dependent weight distribution, enabling information integration and expressive adaptation.
3.2 Weighted Mixture of Experts and Upward Projection
@@
Once we obtain the low-dimensional input $x^{\prime}$, we combine information from all layers using the model-wide shared $B$ experts. One intuitive approach is to feed $x^{\prime}$ into each $B$ expert and combine their outputs to obtain the additional knowledge $x_{\Delta}\in\mathbb{R}^{N\times C_{\text{out}}}$. However, to address efficiency concerns discussed in Appendix A.2, we propose an alternative implementation that is mathematically equivalent but significantly reduces the computational burden, described as follows:
$x_{\Delta}=x^{\prime}\left(\sum_{i=1}^{N_{e}}S_{i}\cdot B^{i}\right)$ (3)
@@
where $S$ is the set of weight scores for the $B$ experts, obtained through selective weight allocation. Since each $S_{i}$ is a scalar value, the calculation in Eq. 3 is mathematically equivalent to the intuitive method but with significantly improved efficiency. Therefore, the complete computation flow, with input $x\in\mathbb{R}^{N\times C_{\text{in}}}$ and output $y\in\mathbb{R}^{N\times C_{\text{out}}}$, for an adaptation target module is:
@@
where $s$ is a scaling factor. By selectively allocating weights and mixing $B$ experts, Lily enables access to all levels of information during adaptation. Each layer’s target adaptation modules can consider the status and knowledge from all other layers, resulting in a more expressive and comprehensive adaptation. Meanwhile, thanks to its interconnectivity, Lily can break the low-rank update constraint of LoRA by simply employing a smaller number of adapters with higher ranks.